Consider a random copolymerization of A and B monomers, where the mole fraction of monomer A is 0.3. What is the mole fraction of B that is being incorporated into the polymer at high monomer conversions?
Answer #1

Since total mole fraction is one .

hence mole fraction of B= 1-0.3= 0.7Polymer = monomer A tronomer B. Since XA + XB = | mole fraction mole fraction of of A B Hence mole fuachon of B Xo=1-03 - 0.7